OpenStudy (anonymous):

A cylindrical grain bin has a radius of 9 feet and a height of 30 feet. How many cubic feet of grain can the bin hold? 1,695.6 cubic feet 7,630.2 cubic feet 847.8 cubic feet 1,907.55 cubic feet

OpenStudy (anonymous):

Volume of the grain bin = area of circle * height = pi * r^2 * h =pi * 9^2 * 30 =7634 cubic feet.